Find the domain of y=x+3if the range is {-1,1,3,5}
I don't get this

Answer by sdmmadam@yahoo.com(530)

Find the domain of y=x+3if the range is {-1,1,3,5}
I don't get this
The range set that is the set of images is {-1,1,3,5}
The correspondence is y = x+3 (which is clearly a 1-1 function that is every image y is having a single preimage x)
so x here is from (x+3) = y
x = y-3 ---(*)
If y = -1 then x = (-1)-3 = -4
Verification: What gets mapped onto (-1) under y = x+3 gives us (-4)as answer and of course (-4) = 3 = -1
Thus if y=-1, then x is -4
Similarly y=1 in (*) gives x = -2
y=3 in (*) gives x = 0
y=5 in (*) gives x = 2
Thus the set that gets mapped onto {-1,1,3,5}
under y = f(x) = x+3 is given by
{-4,-2,0,2}